# Automatic Generation and Evaluation of Reading Comprehension Test Items with Large Language Models

### v1

@@ 4.5.   Automatic (LLM-Based) Evaluation
We used zero-shot prompting to elicit item responses from Llama 2 and GPT-4 in two settings. In the guessing setting, each prompt contained the text, the stem of a single item, and a single answer option, and the models were instructed to respond with a binary (true/false) label. In the comprehension setting, the prompt did not include the text (refer to Appendix A for the full prompts). Both settings used a sampling temperature of 0.
@@
[…]
@@ B.1.   Human-Written Items
German
English
Der Text handelt vor allem von … Fußballfans im Jemen und wie sie versuchen, die Vereine in ihren Orten zu unterstützen. einer großen nationalen Sportveranstaltung, die im Jemen stattfinden sollte. den Vorbereitungen der jemenitischen Nationalmannschaft auf eine wichtige Meisterschaft.
The text is mainly about … Football fans in Yemen and how they try to support the clubs in their towns. a major national sports event that was supposed to take place in Yemen. the preparations of the Yemeni national team for an important championship.
Was sagt Bashir Sinan? Wenn Fußball gespielt wird, vergessen die Jemeniten den Krieg in ihrem Land. Das jemenitische Fußballteam leidet unter den politischen Problemen im Land. Für die Fußballer ist das nächste Ziel, die Asienmeisterschaften zu gewinnen.
What does Bashir Sinan say? When playing football, Yemenis forget about the war in their country. The Yemeni football team suffers from the political problems in the country. The next goal for the football players is to win the Asia Cup.
Was ist richtig? Viele Jemeniten fliehen vor dem Bürgerkrieg in ihrem Land. Die jemenitische Mannschaft will, dass es Friedensgespräche gibt. Katar und Saudi-Arabien kämpfen im Krieg gemeinsam gegen den Jemen.
What is correct? Many Yemenis are fleeing the civil war in their country. The Yemeni team wants peace talks to take place. Qatar and Saudi Arabia are fighting together in the war against Yemen.
@@ B.2.   Items Generated by Llama 2
German
English
Warum dürfen keine Fußballspiele mehr in Jemen stattfinden? Weil das Land zu unsicher ist Weil es ein Bürgerkrieg gibt Weil die Bevölkerung nicht interessiert ist
Why are no more football games allowed in Yemen? Because the country is too unsafe Because there is a civil war Because the population is not interested
Wie haben die jemenitischen Fußballspieler sich für die Asienmeisterschaften qualifiziert? Sie haben in verschiedenen arabischen Ländern trainiert Sie haben ein wildcard [sic] erhalten Sie haben sich nicht qualifiziert
How did the Yemeni football players qualify for the Asia Cup? They trained in various Arab countries They received a wildcard They did not qualify
Was ist die Botschaft der jemenitischen Fußballmannschaft an die Bevölkerung und die Politik? Lassen wir die Gewehre schweigen und einen vernünftigen Dialog und Frieden an ihre Stelle treten Lassen wir die Gewalt weitergehen und uns nichts ausreden Lassen wir uns auf die politischen Spannungen einigen und die Fußballmeisterschaft boykottieren
What is the message of the Yemeni football team to the population and the politicians? Let’s silence the guns and have a reasonable dialogue and peace take their place Let’s let the violence continue and not talk ourselves out of anything Let’s agree on the political tensions and boycott the football championship
@@ B.3.   Items Generated by GPT-4
German
English
Warum dürfen im Jemen keine Fußballspiele mehr stattfinden? Wegen des Bürgerkriegs und der damit verbundenen Unsicherheit. Weil die Fußballstadien zerstört wurden. Weil die jemenitischen Fußballspieler alle das Land verlassen haben.
Why are no more football games allowed in Yemen? Because of the civil war and the associated insecurity. Because the football stadiums were destroyed. Because all Yemeni football players have left the country.
Wo fand das Training für die Asienmeisterschaften statt? In verschiedenen arabischen Ländern, wie Saudi-Arabien und Katar. In Malaysia. Im Jemen.
Where did the training for the Asia Cup take place? In various Arab countries, such as Saudi Arabia and Qatar. In Malaysia. In Yemen.
Was ist die Botschaft der jemenitischen Fußballmannschaft an die Bevölkerung und die Politik? Sie sollten mehr Geld in den Fußball investieren. Sie sollten die Gewehre schweigen lassen und einen vernünftigen Dialog und Frieden an ihre Stelle treten. Sie sollten die Fußballspieler besser unterstützen.
What is the message of the Yemeni football team to the population and the politicians? They should invest more money in football. They should silence the guns and have a reasonable dialogue and peace take their place. They should support the football players better.
@@ D.1.   Guessable Items
Die Saturnalien … wurden an den längsten Tagen im Jahr gefeiert. waren ein Fest, bei dem unfreie Menschen mit den Herrschern die Rollen wechselten. fanden unter anderem in der römischen Stadt Köln statt.
The Saturnalia … were celebrated on the longest days of the year. were a festival where unfree people switched roles with the rulers. took place in the Roman city of Cologne, among other places.
@@
Was bedeutet der Name „Karneval“ aus dem Lateinischen übersetzt? „Fest der Freude“ „Fleisch, leb wohl“ „Tanz der Narren“
What does the name “Carnival” mean when translated from Latin? “Festival of joy” „Flesh, farewell“ „Dance of the fools“
@@ D.2.1.   Wrong label
German
English
Text excerpt
[…] Die Musikwissenschaftlerin Marina Schwarz meint dazu: „Das ist Teil der immer noch patriarchalischen Gesellschaft, in der wir leben.“ Offenbar finden auch viele Frauen, die in dieser Gesellschaft aufgewachsen sind, solche Texte normal. […]
[…] Musicologist Marina Schwarz says: “This is part of the still patriarchal society in which we live.” Apparently, many women who have grown up in this society also find such lyrics normal. […]
@@
Was ist laut Text Marina Schwarz’ Meinung zu sexistischen Texten im Schlager? Sie findet sie inakzeptabel. [should be ✗] Sie findet sie normal, weil es Teil der patriarchalischen Gesellschaft ist. Sie findet sie nicht sexistisch, sondern nur humorvoll.
According to the text, what is Marina Schwarz’ opinion on sexist lyrics in Schlager? She finds them unacceptable. [should be ✗] She finds them normal, because it is part of the patriarchal society. She does not find them sexist, just humorous.
@@ D.2.2.   Unclear answer options
German
English
Text excerpt
[…] Für viele Deutsche zählt beim Kiosk eher die Atmosphäre – besonders in der warmen Jahreszeit. […]
[…] For many Germans, the atmosphere is more important at the kiosk – especially in the warm season. […]
@@
Viele Menschen kaufen Alkohol am Kiosk, weil … er dort billiger ist als in Bars und Kneipen. sie die schöne Stimmung vor Ort mögen. [unclear if ✓or ✗] sie auf dem Weg zu einer Party etwas trinken möchten.
Many people buy alcohol at the kiosk because … it is cheaper there than in bars and pubs. they like the nice atmosphere on site. [unclear if ✓or ✗] they want to drink something on the way to a party.
@@ D.2.3.   Insufficient evidence
German
English
Text excerpt
[…] Besonders im Rheinland sind die Straßen voll mit kostümierten Menschen, die tanzen, singen und feiern […]
[…] Especially in the Rhineland, the streets are full of people in costumes who dance, sing and celebrate […]
@@
Wo finden die meisten Karnevalsumzüge und -feiern statt? In Köln In Rom In Berlin
Where do most carnival parades and celebrations take place? In Cologne In Rome In Berlin

### v2

@@ 4.5.   Automatic (LLM-Based) Evaluation
We used zero-shot prompting to elicit item responses from Llama 2 and GPT-4 in two settings. In the comprehension setting, each prompt contained the text, the stem of a single item, and a single answer option, and the models were instructed to respond with a binary (true/false) label. In the guessing setting, the prompt did not include the text (refer to Appendix A for the full prompts). Both settings used a sampling temperature of 0.
@@
Llama 2 showed a strong bias towards positive responses. We therefore considered the response to be positive if $P(\textit{true})/(P(\textit{true})+P(\textit{false}))\geq\tau$. We optimized the threshold $\tau$ to maximize response accuracy in each setting separately on an additional 50 texts from the same dataset. The resulting thresholds were $\tau_{\text{with text}}=0.9952$ and $\tau_{\text{without text}}=0.9849$. No such optimization was done for GPT-4.
@@ B.1.   Human-Written Items
German English
Der Text handelt vor allem von … Fußballfans im Jemen und wie sie versuchen, die Vereine in ihren Orten zu unterstützen. einer großen nationalen Sportveranstaltung, die im Jemen stattfinden sollte. den Vorbereitungen der jemenitischen Nationalmannschaft auf eine wichtige Meisterschaft. The text is mainly about … Football fans in Yemen and how they try to support the clubs in their towns. a major national sports event that was supposed to take place in Yemen. the preparations of the Yemeni national team for an important championship.
Was sagt Bashir Sinan? Wenn Fußball gespielt wird, vergessen die Jemeniten den Krieg in ihrem Land. Das jemenitische Fußballteam leidet unter den politischen Problemen im Land. Für die Fußballer ist das nächste Ziel, die Asienmeisterschaften zu gewinnen. What does Bashir Sinan say? When playing football, Yemenis forget about the war in their country. The Yemeni football team suffers from the political problems in the country. The next goal for the football players is to win the Asia Cup.
Was ist richtig? Viele Jemeniten fliehen vor dem Bürgerkrieg in ihrem Land. Die jemenitische Mannschaft will, dass es Friedensgespräche gibt. Katar und Saudi-Arabien kämpfen im Krieg gemeinsam gegen den Jemen. What is correct? Many Yemenis are fleeing the civil war in their country. The Yemeni team wants peace talks to take place. Qatar and Saudi Arabia are fighting together in the war against Yemen.
@@ B.2.   Items Generated by Llama 2
German English
Warum dürfen keine Fußballspiele mehr in Jemen stattfinden? Weil das Land zu unsicher ist Weil es ein Bürgerkrieg gibt Weil die Bevölkerung nicht interessiert ist Why are no more football games allowed in Yemen? Because the country is too unsafe Because there is a civil war Because the population is not interested
Wie haben die jemenitischen Fußballspieler sich für die Asienmeisterschaften qualifiziert? Sie haben in verschiedenen arabischen Ländern trainiert Sie haben ein wildcard [sic] erhalten Sie haben sich nicht qualifiziert How did the Yemeni football players qualify for the Asia Cup? They trained in various Arab countries They received a wildcard They did not qualify
Was ist die Botschaft der jemenitischen Fußballmannschaft an die Bevölkerung und die Politik? Lassen wir die Gewehre schweigen und einen vernünftigen Dialog und Frieden an ihre Stelle treten Lassen wir die Gewalt weitergehen und uns nichts ausreden Lassen wir uns auf die politischen Spannungen einigen und die Fußballmeisterschaft boykottieren What is the message of the Yemeni football team to the population and the politicians? Let’s silence the guns and have a reasonable dialogue and peace take their place Let’s let the violence continue and not talk ourselves out of anything Let’s agree on the political tensions and boycott the football championship
@@ B.3.   Items Generated by GPT-4
German English
Warum dürfen im Jemen keine Fußballspiele mehr stattfinden? Wegen des Bürgerkriegs und der damit verbundenen Unsicherheit. Weil die Fußballstadien zerstört wurden. Weil die jemenitischen Fußballspieler alle das Land verlassen haben. Why are no more football games allowed in Yemen? Because of the civil war and the associated insecurity. Because the football stadiums were destroyed. Because all Yemeni football players have left the country.
Wo fand das Training für die Asienmeisterschaften statt? In verschiedenen arabischen Ländern, wie Saudi-Arabien und Katar. In Malaysia. Im Jemen. Where did the training for the Asia Cup take place? In various Arab countries, such as Saudi Arabia and Qatar. In Malaysia. In Yemen.
Was ist die Botschaft der jemenitischen Fußballmannschaft an die Bevölkerung und die Politik? Sie sollten mehr Geld in den Fußball investieren. Sie sollten die Gewehre schweigen lassen und einen vernünftigen Dialog und Frieden an ihre Stelle treten. Sie sollten die Fußballspieler besser unterstützen. What is the message of the Yemeni football team to the population and the politicians? They should invest more money in football. They should silence the guns and have a reasonable dialogue and peace take their place. They should support the football players better.
@@ D.1.   Guessable Items
Die Saturnalien … wurden an den längsten Tagen im Jahr gefeiert. waren ein Fest, bei dem unfreie Menschen mit den Herrschern die Rollen wechselten. fanden unter anderem in der römischen Stadt Köln statt. The Saturnalia … were celebrated on the longest days of the year. were a festival where unfree people switched roles with the rulers. took place in the Roman city of Cologne, among other places.
@@
Was bedeutet der Name „Karneval“ aus dem Lateinischen übersetzt? „Fest der Freude“ „Fleisch, leb wohl“ „Tanz der Narren“ What does the name “Carnival” mean when translated from Latin? “Festival of joy” „Flesh, farewell“ „Dance of the fools“
@@ D.2.1.   Wrong label
German English
Text excerpt […] Die Musikwissenschaftlerin Marina Schwarz meint dazu: „Das ist Teil der immer noch patriarchalischen Gesellschaft, in der wir leben.“ Offenbar finden auch viele Frauen, die in dieser Gesellschaft aufgewachsen sind, solche Texte normal. […] […] Musicologist Marina Schwarz says: “This is part of the still patriarchal society in which we live.” Apparently, many women who have grown up in this society also find such lyrics normal. […]
@@
Was ist laut Text Marina Schwarz’ Meinung zu sexistischen Texten im Schlager? Sie findet sie inakzeptabel. [should be ✗] Sie findet sie normal, weil es Teil der patriarchalischen Gesellschaft ist. Sie findet sie nicht sexistisch, sondern nur humorvoll. According to the text, what is Marina Schwarz’ opinion on sexist lyrics in Schlager? She finds them unacceptable. [should be ✗] She finds them normal, because it is part of the patriarchal society. She does not find them sexist, just humorous.
@@ D.2.2.   Unclear answer options
German English
Text excerpt […] Für viele Deutsche zählt beim Kiosk eher die Atmosphäre – besonders in der warmen Jahreszeit. […] […] For many Germans, the atmosphere is more important at the kiosk – especially in the warm season. […]
@@
Viele Menschen kaufen Alkohol am Kiosk, weil … er dort billiger ist als in Bars und Kneipen. sie die schöne Stimmung vor Ort mögen. [unclear if ✓or ✗] sie auf dem Weg zu einer Party etwas trinken möchten. Many people buy alcohol at the kiosk because … it is cheaper there than in bars and pubs. they like the nice atmosphere on site. [unclear if ✓or ✗] they want to drink something on the way to a party.
@@ D.2.3.   Insufficient evidence
German English
Text excerpt […] Besonders im Rheinland sind die Straßen voll mit kostümierten Menschen, die tanzen, singen und feiern […] […] Especially in the Rhineland, the streets are full of people in costumes who dance, sing and celebrate […]
@@
Wo finden die meisten Karnevalsumzüge und -feiern statt? In Köln In Rom In Berlin Where do most carnival parades and celebrations take place? In Cologne In Rome In Berlin